When copper is heated with an excess of sulfur, copper(I) sulfide is formed. In a given experiment, 1.50 g of copper was heated with excess sulfur to yield 1.53 g copper(I) sulfide. What is the percent yield?


1
Expert's answer
2022-03-10T00:28:01-0500

The equation for the reaction is

2Cu(s) +S(s) "\to" Cu2S(s)

Moles of Cu= 1.5÷63.5 = 0.02362

Expected moles of Cu2S=0.02362÷2

=0.01181

Moles of Cu2S oobtained=1.53÷159

=0.009623

Percentage of yield =(0.009623÷0.01181) ×100 =81.48%